Determine if the following are in proportion.
33, 44, 75, 100

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

33, 44, 75, 100
33/ 44 = 3/ 4
75 / 100 = 3 / 4
Hence, 33:44 = 75: 100
∴These are in a proportion

Thanks